OpenStudy (angel_kitty12):

The lengths of songs on the radio are normally distributed with a mean length of 210 seconds. If 38.2% of all songs have lengths between 194 and 226 seconds, then the standard deviation of this distribution is (1) 16 seconds (2) 32 seconds (3) 8 seconds (4) 64 seconds Can someone please help explain this to me? Thank you!

OpenStudy (lieutenantgeneral):

Do you wanna learn to do it or the answer(even if u pick "answer" ill explain anyways LOL)

OpenStudy (lieutenantgeneral):

38.2% = 382% = 382

OpenStudy (lieutenantgeneral):

16 because its 15.706...... add 194+226+210 = 600 / 38.2 = 15.706.... but if u do 600/38.2% it equals .15706...

OpenStudy (angel_kitty12):

Wait how did you do that? I do want the explanation yea. How did you add those numbers? Where are they from?

OpenStudy (angel_kitty12):

Okay never mind I see the numbers

OpenStudy (angel_kitty12):

Oh alright you divide okay I get it! Thank you so much!